# : Accelerating the Science of Language Models

### v3

@@ Abstract
Language models (LMs) have become ubiquitous in both NLP research and in commercial product offerings. As their commercial importance has surged, the most powerful models have become closed off, gated behind proprietary interfaces, with important details of their training data, architectures, and development undisclosed. Given the importance of these details in scientifically studying these models, including their biases and potential risks, we believe it is essential for the research community to have access to powerful, truly open LMs. To this end, this technical report details the first release of OLMo, a state-of-the-art, truly Open Language Model and its framework to build and study the science of language modeling. Unlike most prior efforts that have only released model weights and inference code, we release OLMo and the whole framework, including training data and training and evaluation code. We hope this release will empower and strengthen the open research community and inspire a new wave of innovation.
@@ 1 Introduction
We believe that full access to open language models for the research community is critical to the scientific study of these models, their strengths and weaknesses, and their biases and risks. Accordingly, we introduce OLMo, a state-of-the-art, truly open language model and framework to build, study, and advance LMs, along with the training data, training and evaluation code, intermediate model checkpoints, and training logs.
Recent LM releases have varied in their degree of openness. For example, Mistral 8x7B provided model weights and a brief report (Jiang et al., 2024), while LLaMA came with in-depth adaptation training instructions (Touvron et al., 2023b), and Mosaic Pretrained Transformer came with many details, including the dataset distribution, though not the data itself (MosaicML NLP Team, 2023). Falcon’s pretraining data was partially released (Almazrouei et al., 2023), and the most open models—the Pythia suite (Biderman et al., 2023) and BLOOM (BigScience et al., 2022)—released training code, model checkpoints, training data and more.
@@
The OLMo framework encompasses the tools and resources required for building and researching language models. For training and modeling, it includes full model weights, training code, training logs, ablations, training metrics in the form of Weights & Biases logs, and inference code. This first release includes four variants of our language model at the 7B scale corresponding to different architectures, optimizers, and training hardware, and one model at the 1B scale, all trained on at least 2T tokens. We are also releasing hundreds of intermediate checkpoints available as revisions on HuggingFace. For dataset building and analysis, it includes the full training data used for these models, including code that produces the training data, from AI2’s Dolma (Soldaini et al., 2024), and WIMBD (Elazar et al., 2023) for analyzing pretraining data. For evaluation, it includes AI2’s Catwalk (Groeneveld et al., 2023) for downstream evaluation and Paloma (Magnusson et al., 2023) for perplexity-based evaluation. For instruction-tuning, we released Open Instruct (Ivison et al., 2023; Wang et al., 2023), and we are currently using it to produce an adapted (instruction-tuned and RLHFed) version of OLMo, which we will release soon. Finally, all code and weights are released under the Apache 2.0 License.111http://www.apache.org/licenses/LICENSE-2.0
This is the first step in a long series of planned releases, continuing with larger models, instruction-tuned models, and more modalities and variants down the line. We therefore hope to catalyze research into as-yet poorly understood aspects of these models, for example, the relationship between pretraining data and model capabilities, the impact of design and hyperparameter choices, and various optimization methods and their impact on model training. In addition, we report on the lessons learned and important details necessary to successfully train language models at this scale.
@@ 2.1 OLMo Model and Architecture
We adopt a decoder-only transformer architecture based on Vaswani et al. (2017), and deliver 1B and 7B variants as described in Table 1, with a 65B version coming soon. Our specific architecture includes several improvements over the vanilla transformer from Vaswani et al. (2017) following other recent large language models like PaLM (Chowdhery et al., 2022), the LLaMA family (Touvron et al., 2023a, b), OpenLM (Gururangan et al., 2023), and Falcon (Almazrouei et al., 2023). Table 2 gives a comprehensive comparison of our 7B architecture to the similarly-sized models from these other families.
Size | Layers | Hidden Size | Attention Heads | Tokens Trained |
1B | 16 | 2048 | 16 | 2T |
7B | 32 | 4086 | 32 | 2.46T |
65B* | 80 | 8192 | 64 |
* At the time of writing our 65B model is still training.
We generally select hyperparameters by optimizing for training throughput on our hardware while minimizing the risk of loss spikes and slow divergence. We ablate choices through our in-loop evaluation setting, given available computational sources (Section 2.4). Table 2 compares our design choices with recent state-of-the-art open language models. Our main changes over the vanilla transformer architecture can be summarized as follows:
@@
Non-parametric layer norm. We use the non-parametric formulation of layer norm (Ba et al., 2016) in which there is no affine transformation within the norm, i.e. no “adaptive gain” (or bias). We believe this was the safest option and it was also the fastest compared to the other variants we considered: parametric layer norm and RMSNorm (Zhang and Sennrich, 2019).
@@
Vocabulary. We use a modified version of the BPE-based tokenizer from GPT-NeoX-20B (Black et al., 2022) with additional tokens for masking personal identifiable information (PII). The final vocabulary size is 50,280. However, to maximize training throughput we increase the size of the corresponding embedding matrix in our model to 50,304 so that it’s a multiple of 128.

### v4

@@ Abstract
Language models (LMs) have become ubiquitous in both NLP research and in commercial product offerings. As their commercial importance has surged, the most powerful models have become closed off, gated behind proprietary interfaces, with important details of their training data, architectures, and development undisclosed. Given the importance of these details in scientifically studying these models, including their biases and potential risks, we believe it is essential for the research community to have access to powerful, truly open LMs. To this end, we have built OLMo, a competitive, truly Open Language Model, to enable the scientific study of language models. Unlike most prior efforts that have only released model weights and inference code, we release OLMo alongside open training data and training and evaluation code. We hope this release will empower the open research community and inspire a new wave of innovation.
@@ 1 Introduction
We believe that full access to open language models for the research community is critical to the scientific study of these models, their strengths and weaknesses, and their biases and risks. Accordingly, we introduce OLMo, a powerful, truly open language model alongside open training data, training and evaluation code, intermediate model checkpoints, and training logs.
Recent LM releases have varied in their degree of openness. For example, Mixtral 8x7B provided model weights and a brief report (Jiang et al., 2024), while LLaMA came with in-depth adaptation training instructions (Touvron et al., 2023b), and Mosaic Pretrained Transformer came with many details, including the dataset distribution, though not the data itself (MosaicML NLP Team, 2023). Falcon’s pretraining data was partially released (Almazrouei et al., 2023), and the most open models—the Pythia suite (Biderman et al., 2023) and BLOOM (BigScience et al., 2022)—released training code, model checkpoints, data, and more.
| Size | L | D | H | Tokens | Peak LR | Warmup | Weight Tying | Batch size |
| 1B | 16 | 2048 | 16 | 2T | 4.0E-4 | 2000 steps | yes | 4M |
| 7B | 32 | 4086 | 32 | 2.46T | 3.0E-4 | 5000 steps | no | 4M |
@@
The OLMo framework encompasses the tools and resources required for building and researching language models. For training and modeling, it includes full model weights, training code, training logs, and inference code. The released model includes four variants of our language model at the 7B scale corresponding to different architectures, optimizers, and training hardware, and one model at the 1B scale, all trained on at least 2T tokens. We also release hundreds of intermediate checkpoints available as revisions on HuggingFace. For dataset building and analysis, the full training data used for these models is openly available (Dolma; Soldaini et al., 2024), including code that produces the training data, and tools for analyzing pretraining data (Elazar et al., 2024). For evaluation, we build on Catwalk (Groeneveld et al., 2023) for downstream evaluation and Paloma (Magnusson et al., 2023) for perplexity-based evaluation. For adaptation, we use Open Instruct (Ivison et al., 2023; Wang et al., 2023) to train with instruction and feedback data. Finally, all code and weights are released under the Apache 2.0 License.111https://allenai.org/olmo
With this release, we hope to catalyze research into as-yet poorly understood aspects of these models, for example, the relationship between pretraining data and model capabilities, the impact of design and hyperparameter choices, and various optimization methods and their impact on model training. In addition, we report on the lessons learned and important details necessary to successfully train language models at this scale.
@@ 2.1 OLMo Model and Architecture
We adopt a decoder-only transformer architecture based on Vaswani et al. (2017), and deliver 1B and 7B variants as described in Table 1. Our specific architecture includes several improvements over the vanilla transformer from Vaswani et al. (2017) following other recent large language models like PaLM (Chowdhery et al., 2022), the LLaMA family (Touvron et al., 2023a, b), OpenLM (Gururangan et al., 2023), and Falcon (Almazrouei et al., 2023). See Table 5 in Appendix A for a comprehensive comparison of our 7B architecture to the similarly-sized models from these other families.
We generally select hyperparameters by optimizing for training throughput on our hardware while minimizing the risk of loss spikes and slow divergence. We ablate choices through our in-loop evaluation setting, given available computational sources (Section 2.4). Our main changes over the vanilla transformer architecture can be summarized as follows:
@@
Non-parametric layer norm. We use the non-parametric formulation of layer norm (Ba et al., 2016) in which there is no affine transformation within the norm, i.e., no “adaptive gain" (or bias). We believe this was the safest option and it was also the fastest compared to the other variants we considered: parametric layer norm and RMSNorm (Zhang and Sennrich, 2019).
@@
Vocabulary. We use a modified version of the BPE-based tokenizer from GPT-NeoX-20B (Black et al., 2022) with additional tokens for masking personal identifiable information (PII). The final vocabulary size is 50,280. However, to maximize training throughput we increase the size of the corresponding embedding matrix in our model to 50,304 to be a multiple of 128.
2.2 Pretraining Data: Dolma
Despite progress in access to model parameters, pretraining datasets are still not as open. Pretraining data are often not released alongside open models (let alone closed models) and documentation about such data is often lacking in detail that would be needed to reproduce or fully understand the work. This has made it difficult to support certain threads of language model research, such as understanding how training data impacts model capabilities and limitations. To facilitate open research on language model pretraining, we built and released our pretraining dataset, Dolma—a diverse, multi-source corpus containing trillions of tokens across billions of documents acquired from different data sources that are (1) commonly seen in large-scale language model pretraining and (2) accessible to the general public (Soldaini et al., 2024). Table 2 provides a high-level overview of the amount of data from each source.
Dolma is built using a pipeline of (1) language filtering, (2) quality filtering, (3) content filtering, (4) deduplication, (5) multi-source mixing, and (6) tokenization. We refer the reader to the Dolma report (Soldaini et al., 2024) for more details about its design principles, details about its construction, and a more detailed summary of its contents. The report provides additional analyses and experimental results from training language models on intermediate states of Dolma to share what we learned about important data curation practices, including the role of content or quality filters, deduplication, and mixing data from multiple sources. We keep documents from each source separate, both during curation as well as in the final release. We open-sourced our high-performance data curation tools; this toolkit can be used to further experiment on Dolma, reproduce our work, and enable fast and easy curation of pretraining corpora. Finally, we also open-sourced our WIMBD tool (Elazar et al., 2024) to help with dataset analysis.
| Source | Type |
| Common Crawl | web pages | 9,812 | 3,734 | 2,180 | |||||||
| GitHub | code | 1,043 | 210 | 342 | |||||||
| social media | 339 | 377 | 80 | ||||||||
| Semantic Scholar | papers | 268 | 38.8 | 57 | |||||||
| Project Gutenberg | books | 20.4 | 0.056 | 5.2 | |||||||
| Wikipedia | encyclopedic | 16.2 | 6.2 | 3.7 | |||||||
| Total | 11,519 | 4,367 | 2,668 | ||||||||
2.3 Adaptation
Pretrained models are not always used as-is, but rather further finetuned to improve their performance, safety, and usability. Often models are first trained to follow instructions (Mishra et al., 2022; Wei et al., 2022; Sanh et al., 2022), and then further trained on human preferences (Ouyang et al., 2022) to improve the quality of their generations. We showcase the efficacy of using OLMo as a base model for further fine-tuning by training OLMo to be a general chat assistant following the Tülu data and training setup (Ivison et al., 2023). This involves first performing instruction finetuning with a mixture of distilled and human-written instruction data and then further aligning the model with distilled preference data using Direct Preference Optimization (DPO) (Rafailov et al., 2023).
2.4 Evaluation
We perform base model evaluation at two stages: online evaluation to make decisions for model design and offline evaluation to evaluate model checkpoints. For the offline stage, we use the Catwalk framework (Groeneveld et al., 2023), a publicly available evaluation tool with access to a wide range of datasets and task formats, to perform downstream evaluation as well as intrinsic language modeling evaluation on the perplexity benchmark Paloma (Magnusson et al., 2023).
In-Loop Training Ablations
Downstream Evaluation
Intrinsic Language Modeling Evaluation
To measure how OLMo fits distributions of language beyond held-out training data, we use Paloma (Magnusson et al., 2023), a new perplexity benchmark that includes 585 different domains of text. Domains range from nytimes.com to r/depression on Reddit and are drawn from 18 separate data sources, such as C4 (Raffel et al., 2020), in stratified samples. This allows for more equal inclusion of text domains that are under-represented in their source corpora.
We aim not just to compare OLMo against other models for best performance, but also to demonstrate how it enables fuller and more controlled scientific evaluations. OLMo-7B is the largest LM with explicit decontamination for perplexity evaluation. Following the approach described in Paloma, we remove any pretraining document with paragraphs leaked from Paloma evaluation data. Without decontamination, other models risk underestimating perplexity (i.e., overestimating the model’s out-of-sample fit). We also release intermediate checkpoints, allowing richer comparisons with two other models that release checkpoints, Pythia-6.9B (Biderman et al., 2023) and RPJ-INCITE-7B (Together Computer, 2023) (see Figure 2).
Adaptation Evaluation
We also evaluate OLMo after instruction fine-tuning and DPO training using the Tülu evaluation suite proposed in Wang et al. (2023); Ivison et al. (2023). We focus on evaluations around model chat capabilities and safety in order to showcase the efficacy of using OLMo as a base for further fine-tuning.
3 Training OLMo
This section describes our pretraining setup, including our distributed training framework (Section 3.1), optimizer (Section 3.2), data preparation (Section 3.3), and hardware (Section 3.4).
3.1 Distributed Training Framework
We train our models using the ZeRO optimizer strategy (Rajbhandari et al., 2019) via PyTorch’s FSDP framework (Zhao et al., 2023), which reduces memory consumption by sharding the model weights and their corresponding optimizer state across GPUs. At the 7B scale, this enables training with a micro-batch size of 4096 tokens per GPU on our hardware (see Section 3.4). For OLMo-1B and -7B models, we use a constant global batch size of approximately 4M tokens (2048 instances, each with a sequence length of 2048 tokens).
3.2 Optimizer
3.3 Data
3.4 Hardware
4 Results
The checkpoint used for evaluating OLMo-7B is trained until 2.46T tokens on the Dolma (Soldaini et al., 2024) dataset with a linear learning rate decay schedule mentioned in Section 3.2. In our experiments, we find that tuning this checkpoint further on the Dolma dataset for 1000 steps with the learning rate linearly decayed to 0 boosts model performance on perplexity and end-task evaluation suites described in Section 2.4. We compare OLMo with other publicly available models including LLaMA-7B (Touvron et al., 2023a), Llama-2-7B (Touvron et al., 2023b), MPT-7B (MosaicML NLP Team, 2023), Pythia-6.9B (Biderman et al., 2023), Falcon-7B (Almazrouei et al., 2023) and RPJ-INCITE-7B (Together Computer, 2023).
| Models |
boolq |
piqa | sciq |
avg. | ||||||||||
| StableLM 1.6B | 43.8 | 63.7 | 76.6 | 68.2 | 45.8 | 74.0 | 94.7 | 64.9 | 66.5 | ||||||||||
| Pythia 1B | 33.1 | 50.2 | 61.8 | 44.7 | 37.8 | 69.1 | 86.0 | 53.3 | 54.5 | ||||||||||
| TinyLlama 1.1B | 34.8 | 53.2 | 64.6 | 58.7 | 43.6 | 71.1 | 90.5 | 58.9 | 59.4 | ||||||||||
| [] OLMo-1B | 34.5 | 58.1 | 60.7 | 62.5 | 46.4 | 73.7 | 88.1 | 58.9 | 60.4 | ||||||||||
| Falcon-7B | 47.5 | 70.4 | 74.6 | 75.9 | 53.0 | 78.5 | 93.9 | 68.9 | 70.3 | ||||||||||
| LLaMA 7B | 44.5 | 67.9 | 75.4 | 76.2 | 51.2 | 77.2 | 93.9 | 70.5 | 69.6 | ||||||||||
| Llama 2 7B | 48.5 | 69.5 | 80.2 | 76.8 | 48.4 | 76.7 | 94.5 | 69.4 | 70.5 | ||||||||||
| MPT-7B | 46.5 | 70.5 | 74.2 | 77.6 | 48.6 | 77.3 | 93.7 | 69.9 | 69.8 | ||||||||||
| Pythia 6.9B | 44.1 | 61.9 | 61.1 | 63.8 | 45.0 | 75.1 | 91.1 | 62.0 | 63.0 | ||||||||||
| RPJ-INCITE-7B | 42.8 | 68.4 | 68.6 | 70.3 | 49.4 | 76.0 | 92.9 | 64.7 | 66.6 | ||||||||||
| [] OLMo-7B | 48.5 | 65.4 | 73.4 | 76.4 | 50.4 | 78.4 | 93.8 | 67.9 | 69.3 |
4.1 Downstream evaluation
Setup
Our core downstream evaluation suite (see Table 3) consists of: arc (both arc_easy and arc_challenge) (Clark et al., 2018), boolq (Clark et al., 2019), openbookqa (Mihaylov et al., 2018), sciq (Welbl et al., 2017), hellaswag (Zellers et al., 2019), piqa (Bisk et al., 2020), and winogrande (Sakaguchi et al., 2021). In Appendix C, we also report results on an additional set of auxiliary tasks outside of our core evaluation set that we found to have less stable performance trends (see Figure 4).
In all cases, we perform zero-shot evaluation using the rank classification approach popularized by Brown et al. (2020). Under this approach, candidate text completions (e.g., different multiple-choice options) are ranked by likelihood (usually normalized by some normalization factor), and prediction accuracy is reported. While Catwalk implements several common likelihood normalization strategies, including normalizing by number of tokens (per-token normalization; Brown et al., 2020; Liang et al., 2022), by number of characters (per-character normalization; Gao et al., 2023), as well as incorporating an answer’s unconditional likelihood (Brown et al., 2020), we selected the normalization strategies for each dataset separately. Specifically, we used unconditional normalization for arc and openbookqa, per-token normalization for hellaswag, piqa, and winogrande and no normalization for boolq, and sciq (i.e., tasks formulated as single token prediction tasks).
Results
Table 3 summarizes the result of zero-shot evaluation of OLMo and compares against other publicly available models of comparable size. We report results on 8 core tasks from our evaluation suite described in Section 2.4. On aggregate, OLMo-7B is competitive against all the comparable models. We include the comparison to StableLM 1.6B , but note that it is significantly larger, and was trained on unknown data.
In Figure 1 we plot the accuracy score progression of 8 core end-tasks. All tasks, except OBQA, show an upward trend in accuracy numbers as OLMo-7B is trained on more tokens. A sharp upward tick in accuracy of many tasks between the last and the second to last step shows us the benefit of linearly reducing the LR to 0 over the final 1000 training steps. See Table 7 in Appendix C for additional evaluation results and discussion.
4.2 Intrinsic language modeling evaluation
Setup
Results
The remaining subplots in Figure 2 provide more fine-grained analysis by reporting bits per byte separately for each of the 11 data sources that are combined in the aggregated Paloma metric. From this we see greater variation in sample efficiency, largely driven by the similarity of training and evaluation distributions. Notably, OLMo-7B fares well on evaluations predominated by Common Crawl, such as C4, though different ways of postprocessing Common Crawl are best fit by models trained with that specific data, such as Falcon-7B on Falcon RefinedWeb. Meanwhile, OLMo-7B is less sample efficient compared to other models on sources less related to scraped web text, such as WikiText-103, M2D2 S2ORC, and M2D2 Wikipedia. The RedPajama evaluation shows a similar pattern, perhaps as only 2 of its 7 domains are from Common Crawl, and Paloma weights domains within each source equally. Since heterogeneous data from curated sources like Wikipedia and ArXiv papers is scarcer than scraped web text, maintaining sample efficiency for fit to these distributions of language will be challenging as pretraining corpora are scaled.
4.3 Adaptation Evaluation
Model MMLU AlpacaEval ToxiGen TruthfulQA 0-shot %win % Toxic %Info+True OLMo (base) 28.3 - 81.4 31.6 MPT Chat 33.8 46.8 0.1 42.7 Falcon Instruct 25.2 14.0 70.7 27.2 RPJ-INCITE Chat 27.0 38.0 46.4 53.0 Llama-2-Chat 46.8 87.3 0.0 26.3 Tülu 2 50.4 73.9 7.0 51.7 Tülu 2+DPO 50.7 85.1 0.5 -777Following Ivison et al. (2023), we do not report Tülu 2 TruthfulQA scores due to test set contamination. [] OLMo+SFT 47.3 57.0 14.4 41.2 [] OLMo+SFT+DPO 46.2 69.3 1.7 52.0
Setup
We evaluate OLMo-7B before adaptation, and after both the supervised fine-tuning and DPO training stage, focusing on the safety and chat evaluations used by Wang et al. (2023). We additionally compare to officially released instruction-tuned variants of the models from Table 3. We finally also compare to Tülu 2 models to compare against models trained using the same post-training data mixes and procedures.
Results
We find that instruction tuning considerably improves the performance and safety of OLMo-7B, increasing MMLU performance by a wide margin and improving ToxiGen and TruthfulQA scores - especially after DPO training. Additionally, we find that OLMo-7B outperforms most other chat variants after both initial instruction tuning (OLMo+SFT) and additional preference alignment (OLMo+SFT+DPO), highlighting both the strength of OLMo-7B as a base model and the strength of the Tülu mix used to perform adaptation training. However, we find there is still a gap with Tülu 2, which is trained by applying the Tülu mix on Llama 2. This gap may be due to test set contamination in Llama 2888Touvron et al. (2023b) report that Llama 2 was pretrained on data contaminated with MMLU test data. and because the Tülu mix was primarily designed for Llama models. Overall, we see that OLMo-7B greatly benefits from additional tuning and serves as a strong base model for downstream applications.
5 Artifacts Released
Pretraining (§2.1)
1.
The training and modeling code.
2.
The trained model weights for the 7B model, 7B-twin-2T, and the 1B model. For all the models, we release not only the final model weights but also 500+ intermediate checkpoints at intervals of 1000 steps.
3.
The complete set of metrics logged to Weights & Biases during training.
1.
Adaptation (§2.3)
1.
The training code and data for adaptation.
2.
The model weights for OLMo+SFT and OLMo+SFT+DPO.
1.
6 Conclusion and Future Work
This paper presents our first release of OLMo, a state-of-the-art, truly open language model and its framework to build and study the science of language modeling. Unlike most prior efforts that have only released model weights and inference code, we release OLMo and the whole framework, including training data, training and evaluation code, and detailed metrics collected during the training runs. Additionally, we released adapted models, as well as all of our model adaptation code and data.
We intend to continuously support and extend OLMo and its framework, and continue to push the boundaries of open LMs to empower the open research community. Since the original release of OLMo described here, we improved our data and training setup to significantly improve results. For example, MMLU scores have improved by 24 points to 52%.999 https://medium.com/p/92b43f7d269d We look forward to bringing different model sizes, modalities, datasets, safety measures, and evaluations into the OLMo family. We hope this and future releases will empower and strengthen the open research community and inspire a new wave of innovation.
Limitations
We recognize building a large language model has many limitations. In fact, each step of the process of creating a language model, from the data to training to adaptation to evaluation each have their own limitations, and so we’ve added sections for each below. Of course we recognize that AI systems today can have broad societal reach, and therefore there are significant limitations beyond what we are able to fit into this section.
Data
Our work focuses on pretraining data in English. We hope that our open framework enables the development of future models in more languages as well as multilingual models. The data that models are trained on is what gives models their capabilities, and at the scale of training a large language model we recognize that the data likely contains problematic content like toxic language, personal information, and copyrighted text. We mitigated this to the best of our ability but recognize there are no perfect approaches today that can completely remove such content.
Training
Training a large language model is currently a challenging endeavor which is missing significant support from the open source community. With our limited page count we did not provide extensive training logs documenting, for example, training runs that diverged or failed to learn.
Adaptation
Our pretrained models face the same issues as existing pretrained LLMs, such as bias, toxicity and, hallucinations. Our adapted models are better at avoiding these generations, but they are not perfect. Additionally, we note that we largely adopt an existing data mixture designed for a different model family (Tülu, designed for Llama models), and OLMo may require different data mixing to adjust for its unique strengths and weaknesses. The Tülu mix itself also relies on data distilled from a variety of models, and we hope to reduce our reliance on such data in the future.
Evaluation
While we’ve included comparisons on a variety of datasets to other current language models, many of the downstream tasks are not actually representative of how users interact with language models (i.e., as a chatbot). In addition, language model evaluations are currently very noisy; we aimed to include only evaluations on datasets that provided some signal as to which model performs best, but recognize that there is no perfect automatic evaluation, and thus comparisons should be taken with a grain of salt.
Ethics Statement
Through this work, we take the position that increased openness of language models is essential for scientific understanding of their abilities and limitations and for broad participation in the continued development of such models. Training on open data further enhances these benefits. In addition, our open release enables practitioners to take our models and build on them instead of having to train their own from scratch, in which case they would be repeating our work while consuming more resources and leading to an increased environmental impact. Of course, openness is not without risk; the possibility remains that these models will be used in unintended ways that cause harm. We believe that research and development efforts to understand and mitigate those potential harms will also be accelerated by the openness of the models, allowing a diversity of approaches and analyses. Over the past year there have been a number of comparable models released with very permissive licenses, so using a more strict license for our work would not remove the overall risk in the field. We believe this trade-off on the side of being more open is the best option.
Acknowledgments
Appendix A Training Settings
Table 5 summarizes the model architecture and the optimizer parameters of OLMo-7B as well as recent similar-sized models.